P and Q are two stations. Train A started from P
towards Q at 6:00 a.m at 90 kmph. At the same
time, train B started from R, an intermediate station
60 km from P, and travelled towards Q at 60 kmph.
Train C started from Q towards P at 7:00 a.m at
120 kmph. All the trains crossed each other
simultaneously. Find PQ (in km)

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52812)

.
P and Q are two stations. Train A started from P
towards Q at 6:00 a.m at 90 kmph. At the same
time, train B started from R, an intermediate station
60 km from P, and travelled towards Q at 60 kmph.
Train C started from Q towards P at 7:00 a.m at
120 kmph. All the trains crossed each other
simultaneously. Find PQ (in km)
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Let t be the time train A traveled to get the meeting point (in hours).


     Then from the problem, train B traveled the same time t from R to get the meeting point;
                            train C traveled (t-1) hours from Q to get the meeting point.

 
Consider trains A and B.

     To get the meeting point, train A traveled 90*t kilometers;
                               train B traveled 60*t kilometers.

     For train B, the travel distance to get the meeting point was 60 km less than that for train A, 
     so we write this equation

          90t = 60t + 60.


We simplify this equation and find t

          90t - 60t = 60

           30t = 60

             t = 60/30 = 2 hours.


Now we know that the travel time for A to get the meeting point was 2 hours;
hence, the travel time for train C was 1 hour less that that, or 2-1 = 1 hour.


At this point, we can find the total distance PQ.  It is the sum of partial distances 
for train A and train C to get the meeting point, or

    90*2 + 120*1 = 180 + 120 = 300 kilometers.


ANSWER.  The distance from P to Q is 300 kilometers.

Solved.

------------------------

As you see from this solution, the basic skills required is an ability to solve
standard basic Travel @ Distance problems, and combining different aspects of them.

For simple Travel & Distance problems,  see introductory lessons
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many basic problems there.

Read them and learn once and for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.



Answer by greenestamps(13203)



Trains A and B both started at the same time going in the same direction at different speeds, with B having a "head start" of 60km.

Train A catches up to train B at a rate equal to the difference of their speeds -- at 30km each hour.

Since train A catches up to train B at a rate of 30km each hour, and since it needs to make up a distance of 60km, the number of hours it takes train A to catch up to train B is 60/30 = 2.

So when train A catches up to train B, it has traveled 2 hours at 90km/h, a distance of 180km; train B started 60km from P and traveled 2 hours at 60km/h, putting it at a distance of 60+2(60) = 180km from P.

Train C started from Q and headed towards P at 120km/h. I started 1 hour later that the other two trains, so it traveled 1 hour at 120km/h, a distance of 120km.

Trains A, B, and C reached the same place at the same time. At that time, trains A and B were 180km from P and train C was 120km from Q. So the distance from P to Q was 180+120 = 300km.

ANSWER: 300km